# Fooling the Textual Fooler via Randomizing Latent Representations

### v1

@@ 4.1 Experimental setup
BERT-base (baseline) 94.20 35.20 (62.63) 351.92 36.80 (60.93) 317.27 46.50 (50.64) 337.84
@@
Evaluation Metrics. We evaluate the performance of AdvFooler using four evaluation metrics: Clean accuracy (Clean%), Accuracy under attack (AuA%), Attack success rate (ASR%), Average number of queries (#Query). Cleanmeasures the impact of a defense method on the clean performance of the model. Both AuAand ASRmeasure the robustness of the model under attack. #Query measures how many times, on average, the adversary has to query the model to find adversarial examples; the higher #Query is, the more difficult the attack is.
Datasets. We evaluate AdvFooler on two benchmark datasets: AG-News Corpus (AGNEWS(Zhang et al., 2015) and Internet Movie Database (IMDB(Maas et al., 2011). AGNEWS is a classification dataset of news articles created from the AG’s corpus [^1] http://groups.di.unipi.it/~gulli/AG_corpus_of_news_articles.html. It contains 120,000 training samples and 7,600 test samples where a sample belongs to one of four classes: World, Sports, Business, and Sci/Tech. IMDB is a binary sentiment classification dataset of reviews extracted from the IMDB website. This dataset contains 50,000 samples and is split equally into the training and test sets.
Model Architecture. To assess the generality of AdvFooler on model architectures, we evaluate it on two state-of-the-art, pre-trained language models: BERTbase (Devlin et al., 2019) and RoBERTabase (Liu et al., 2019). BERT is a Transformer-based language model that outperforms other models in various benchmarks at the time of its release and is widely used for text classification tasks. RoBERTa is a variant of BERT with an improved training process and performance.
Adversarial Attacks and Parameters. We choose the following widely used and state-of-the-art adversarial attacks to evaluate our defense: TextFooler (Jin et al., 2020), TextBugger (Li et al., 2018), and BERT-Attack (Li et al., 2020). TextFooler utilizes the word embeddings from (Mrkšić et al., 2016) to generate the set of synonyms for each word, uses both part-of-speech and semantic similarity, and selects words for replacements using important scores. BERT-Attack is similar to TextFooler but uses a pretrained BERT to generate the set of synonyms for a word, given the context of the sentence. TextBugger locates the tokens for perturbations by calculating the important scores for each sentence in the input and then calculating the important score of each token in the sentence. In addition to a synonym set for word substitution, TextBugger also introduces word augmentation - perturbing the characters in words. For a fair evaluation, the adversarial attacks used in the evaluation will be based on Li et al. (2021b), which is commonly used by other papers for benchmarking adversarial defenses (Nguyen Minh & Luu, 2022; Zhang et al., 2022)
@@
Baseline Defenses. We compare AdvFooler against various types of defense methods. For empirical defenses, we select Adversarial Sparse Convex Combination (ASCC) (Dong et al., 2021), InfoBERT (Wang et al., 2020), FreeLB (Zhu et al., 2020), and Textual Manifold Defense (TMD) (Nguyen Minh & Luu, 2022). ASCC, InfoBERT, and FreeLB improve the model’s robustness either by employing adversarial training or introducing additional regularization to the model’s training process. ASCC models the word substitution attack space as a convex hull and leverages a regularization term to generate adversarial examples. Different from ASCC, FreeLB introduces adversarial perturbations to word embeddings and minimizes the corresponding adversarial loss around input samples. InfoBERT adds two regularization terms to eliminate noisy information between the input and hidden features and to increase the correlations between the local and global features. Textual Manifold Defense introduces a defense mechanism that trains a mapping function to project the embeddings of input sentences to a set of predefined manifolds to reduce the effects of the adversarial examples. For certified defenses, RanMASK (Zeng et al., 2021) and SAFER (Ye et al., 2020) are selected. Given an input, both RanMASK and SAFER construct a set of randomly perturbed inputs, using synonym substitution for SAFER and token masking for RanMASK, respectively. Then, they leverage the statistical properties of the predicted output to achieve better model robustness.
@@
Implementation Details. We follow the implementation guidelines for evaluating adversarial attacks and defenses in Li et al. (2021b): (1) the maximum percentage of modified words $\rho_{max}$ for AGNEWS and IMDB are 0.3 and 0.1, respectively, (2) the maximum number of candidate replacement words $K_{max}$ is set to 50, (3) the maximum percentage of modified words $\rho_{max}$ for AGNEWS and IMDB must be 0.3 and 0.1, respectively, and (4) the maximum number of queries to the victim model is $Q_{max}=K_{max}*L$, where $L$ is the length of, or the number of tokens in, the input sentence. Cleanis calculated on the entire test set from each dataset, while the robustness metrics AuA%, ASR%, and #Query are computed on 1,000 randomly chosen samples from the test set. All attacks in the experiments are based on the implementation of TextAttack (Morris et al., 2020).
4.2 Defense performance against textual adversarial attacks
@@
We report the defense performance on AGNEWS in Table 2. As we can observe in this table, on the base model BERT, AdvFooler outperforms all adversarial training techniques (ASCC, FreeLB, and InfoBERT) and the randomized smoothing method, SAFER. The robustness of AdvFooler is comparable to TMD’s robustness and slightly lower than the robustness of another randomized-smoothing defense, RanMASK. However, the clean accuracy in RanMASK drops significantly, approximately 4%, while the clean accuracy in AdvFooler is guaranteed to be within 1% of the base model. Note that, all of the evaluated textual defenses require access to the training data. On the base model RoBERTa, AdvFooler achieves similar robustness to the best defenses, but without the additional computational overhead (see supplementary material for the computational overhead experiment) or access to the training data.
@@
For IMDB’s results in Table 3, we can observe similar robustness results. AdvFooler is effective against the attacks and achieves comparable results to the best defenses (SAFER and TMD).
@@
Difficulty of Adversarial Attacks against AdvFooler. For all the considered adversarial attacks, the number of queries to locate important words (the first step) of an input text is the same with or without the defenses. Thus, the higher number of queries in an experiment is a result of a more difficult synonym-replacement phase. This also explains considerably smaller average numbers of queries for the baseline (i.e., without any defense) in Tables 2 and 3, compared to the protected models.
@@
AdvFooler with other attacks. In the supplementary materials, we provide additional experiments evaluating AdvFooler  against other types of attacks, including black-box hard-label (Appendix LABEL:sec:hard-label) and white-box attacks (Appendix LABEL:sec:white-box), even though they do not align with the threat model studied in this work and several related works (Nguyen Minh & Luu, 2022; Zhang et al., 2022; Zeng et al., 2021; Li et al., 2023) to AdvFooler. The results show that AdvFooler  is still effective in defending against these attacks, further highlighting the versatility of AdvFooler.
@@ 4.3 Robustness from randomizing different latent spaces
In this section, we study the model’s robustness when randomizing different layers of the model. We also decided to study the effects of adding noise to only one layer of the model. For comparison, Specifically, we randomize only the first, middle, and last layers and all layers of the model. The model’s robustness against TextFooler is reported in Figure 3. As we can observe from this figure, randomizing the early layer of the model is more effective than randomizing the later layers of the model. More importantly, adding noise to all the layers leads to a significantly higher accuracy under attack and a higher average number of queries.

### v2

@@ 3.4 Empirical analysis
BERT-base (baseline) 95.14 35.20 (62.63) 351.92 36.80 (60.93) 317.27 46.50 (50.64) 337.84
@@
4 Experiments
4.1 Experimental setup
Evaluation Metrics. We evaluate the defenses using four metrics: Clean accuracy (Clean%), Accuracy under attack (AuA%), Attack success rate (ASR%), Average number of queries (#Query). Cleanmeasures the defense’s impact on the clean performance. Both AuAand ASRmeasure the model’s robustness under attack. #Query is the averaged number of queries used by the adversary to find adversarial examples; the higher #Query is, the more difficult the attack is.
Datasets and Model Architectures. We evaluate the defenses on two benchmark datasets: AG-News Corpus (AGNEWS(Zhang et al., 2015) and Internet Movie Database (IMDB(Maas et al., 2011). To assess the generality of AdvFooler on model architectures, we evaluate it on two state-of-the-art, pre-trained language models: BERTbase (Devlin et al., 2019) and RoBERTabase (Liu et al., 2020). More details are provided in the Appendix.
Adversarial Attacks and Parameters. We choose the following widely used and state-of-the-art adversarial attacks to evaluate our defense: TextFooler (Jin et al., 2020), TextBugger (Li et al., 2018), and BERT-Attack (Li et al., 2020). Both TextFooler and BERT-Attack rely on the important scores to identify important words and replace them with the corresponding synonyms; TextFooler uses word embeddings from (Mrkšić et al., 2016) for synonym generations, while BERT-Attack uses pretrained BERT to generate the synonyms. In addition to replacing words, TextBugger also introduces word augmentation - perturbing the characters in words.
@@ 4.1 Experimental setup
Baseline Defenses. We compare AdvFooler against various types of defense methods. For empirical defenses, we select Adversarial Sparse Convex Combination (ASCC) (Dong et al., 2021), InfoBERT (Wang et al., 2021), FreeLB (Zhu et al., 2020), and Textual Manifold Defense (TMD) (Nguyen Minh and Luu, 2022). ASCC and FreeLB improve the model’s robustness by employing adversarial training; FreeLB introduces adversarial perturbations to word embeddings, while ASCC generates adversarial examples from a convex hull spanned by word substitution. InfoBERT incorporates information theory to refine both local and global features. Textual Manifold Defense reduces the effects of the adversarial examples by projecting the embeddings of input sentences to an approximated embedding manifold. For certified defenses, we evaluate on RanMASK (Zeng et al., 2023) and SAFER (Ye et al., 2020), which construct a set of randomly perturbed inputs using token masking and synonym substitution, respectively. Then, they leverage the statistical properties of the predicted output to achieve better model robustness.
@@
Implementation Details. For a fair evaluation, we follow the implementation guidelines for evaluating adversarial attacks and defenses in Li et al. (2021b): (1) the maximum percentage of modified words $\rho_{max}$ for AGNEWS and IMDB are 0.3 and 0.1, respectively, (2) the maximum number of candidate replacement words $K_{max}$ is set to 50, (3) the maximum percentage of modified words $\rho_{max}$ for AGNEWS and IMDB must be 0.3 and 0.1, respectively, and (4) the maximum number of queries to the victim model is $Q_{max}=K_{max}*L$, where $L$ is the length of, or the number of tokens in, the input sentence. Cleanis calculated on the entire test set from each dataset, while the robustness metrics AuA%, ASR%, and #Query are computed on 1,000 randomly chosen samples from the test set. All attacks are based on TextAttack implementation (Morris et al., 2020).
4.2 Defense performance
@@
We report the defense performance on AGNEWS in Table 2. As we can observe in this table, on the base model BERT, AdvFooler outperforms all adversarial training techniques (ASCC, FreeLB, and InfoBERT) and the randomized smoothing method, SAFER. AdvFooler is consistently in the top-3 robust defenses; the robustness of AdvFooler and TMD are within 1-2% of each other, while they are generally both lower than the robustness of another randomized-smoothing defense, RanMASK. However, the clean accuracy in RanMASK drops significantly, approximately 4%, while the clean accuracy in AdvFooler is guaranteed to be within 1% of the base model. Note that, all of the evaluated textual defenses require access to the training data. On the base model RoBERTa, AdvFooler similarly achieves a top-3 performance.
@@
For IMDB’s results in Table 3, we can observe similar robustness results. TMD’s performance is observed to be consistently better with RoBERTa-base. AdvFooler is effective against the attacks and achieves comparable results to the best defenses (SAFER and TMD).
In summary, AdvFooler achieves competitive robustness to the state-of-the-art defenses; however, AdvFooler has significantly lower training and inference overhead compared to TMD, RandMASK, and SAFER and does not have access to the training data.
@@
Difficulty of Adversarial Attacks against AdvFooler. For all the considered adversarial attacks, the number of queries to locate important words (the first step) of an input text is the same with or without the defenses. Thus, the higher number of queries in an experiment is a result of a more difficult synonym-replacement phase. This also explains considerably smaller average numbers of queries for the baseline (without any defense) in Tables 2 and 3, compared to the protected models.
@@
AdvFooler with other attacks. In the Supplementary, we provide additional experiments evaluating AdvFooler  against other types of attacks, including black-box hard-label (Appendix 4.7) and white-box attacks (Appendix C.5), even though they do not align with the threat model studied in this work and several related works (Nguyen Minh and Luu, 2022; Zhang et al., 2022; Zeng et al., 2023; Li et al., 2023) to AdvFooler. The results show that AdvFooler  is still effective in defending against these attacks, further highlighting the versatility of AdvFooler.
@@ 4.3 Robustness from randomizing different latent spaces
In this section, we study the model’s robustness when randomizing different layers of the model. For comparison, Specifically, we randomize only the first, middle, and last layers and all layers of the model. The model’s robustness against TextFooler is reported in Figure 3. As we can observe, randomizing the early layer of the model is more effective than randomizing the later layers of the model. More importantly, adding noise to all the layers leads to a significantly higher accuracy under attack and a higher average number of queries.
@@ 4.5 AdvFooler with adversarially trained models
4.6 Inference time analysis
Table 5: Inference time comparison with other defenses. Tested with BERT on the test set from the IMDB dataset.
Defense Runtime (s) % Increase
Baseline 75
ASCC 90 ($+$20.0%)
FreeLB 75 ($+$0.0%)
InfoBERT 75 ($+$0.0%)
TMD 100 ($+$33.3%)
RanMask 1752 ($+$2236.0%)
SAFER 1772 ($+$2262.6%)
AdvFooler 77 ($+$2.6%)
In this experiment, we study the computational overhead that each defense introduces at inference time. We use the IMDB test set and record the inference runtime of the BERT model under different defenses. The results are shown in Table 5. From the result, we can see that AdvFooler introduces very little computational overhead compared to most other defenses. Randomized-smoothing methods, such as RanMask and SAFER, incur significant overhead to the inference phase. TMD also adds non-trivial inference overhead (33.3%); as acknowledged by the authors, the main bottleneck of TMD is its high computational overhead of the on-manifold projection (Nguyen Minh and Luu, 2022). Only adversarial training methods, such as ASCC and FreeLB, do not introduce additional overhead during inference time. Most importantly, AdvFooler is the only method that has consistently high robustness and a negligible inference overhead (2.6%).
4.7 AdvFooler’s performance on hard-label attacks.
Table 6: Hard-label results against base models with and without AdvFooler under different noise scale.
Results [AuA%(ASR%)]
BERT-base 58.3% (38.11%)
AdvFooler (0.7 scale) 78.8% (15.35%)
AdvFooler (0.9 scale) 83.3% (9.45%)
AdvFooler (1.1 scale) 85.7% (5.72%)
In some cases, the adversary can choose to ignore the output logits and only rely on the hard-label output of the model, as seen in Maheshwary et al. (2021). The so-called hard-label attacks first initialize a sample that flips the prediction, move it toward the original input until it passes through the decision boundary, then select the sample at the previous step as adversarial example. However, since AdvFooler is a randomized model, it might sometimes return labels that are different from the model without noise for the attack queries. Therefore, it can fool the attack to go a bit further, and the sample found by the attack turns out to be non-adversarial. In this section, we study the performance of AdvFooler against this hard-label attack. In this experiment, we ran these attacks against the base model and AdvFooler. The attack may accidentally flag a chosen perturbation as an adversarial example although this sample may be mispredicted due to the added noise in our defense; in other words, the selected sample is not a true adversarial example since, under a different random noise, the model prediction may still be correct. To avoid such a situation and fairly evaluate the performance of AdvFooler , for each “supposedly adversarial example” generated by the attack, we feed it to the model 5 times and record the model’s predictions; the model’s prediction used for performance evaluation is the prediction occurred most. As we can observe from Table 6, hard-label attack struggles to find consistent Adversarial Examples the bigger the noise being inserted into the model.